A particle moves from x1 = 5 cm to x2 = 20 cm in 5 s. The average velocity of the particle during this time interval is


1
Expert's answer
2021-11-14T17:18:20-0500
"v_{avg}=\dfrac{\Delta x}{\Delta t}=\dfrac{0.2\ m-0.05\ m}{5\ s}=0.03\ \dfrac{m}{s}=3\ \dfrac{cm}{s}."